OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = 2x2 – 8x + 6? Answer x–intercepts: (–3, 0) and (–1, 0) y–intercept: (0, 6) x–intercepts: (–3, 0) and (–1, 0) y–intercept: (0, 3) x–intercepts: (–3, 0) and (–1, 0) y–intercept: (0, 3) x–intercepts: (3, 0) and (1, 0) y–intercept: (0, 6)

OpenStudy (anonymous):

x-intercepts occur where y = 0. 0 = 2x² - 8x + 6 0 = 2(x² - 4x + 3) 0 = 2(x - 3)(x - 1) x - 3 = 0 and x - 1 = 0 x = {3,1}

OpenStudy (anonymous):

thanks so much =]